\begin{thebibliography}{99}

\bibitem{Amghibech}
S. Amghibech, \emph{On the Borel-Cantelli Lemma and moments},
Comment. Math. Univ. Carolin.  47  (2006) 669--679.

\bibitem{Dawson}
D.~A. Dawson, D. Sankoff, \emph{An inequality for probabilities},
Proc. Amer. Math. Soc.  18 (1967) 504--507.

\bibitem{ErdosRenyi}
P. Erd\H{o}s, A. R\'{e}nyi, \emph{On Cantor's series with convergent
$\sum1/q_n$}, Ann. Univ. Sci. Budapest. E\"{o}tv\"{o}s Sect. Math. 2
(1959) 93--109.


\bibitem{Gallot}

S. Gallot, \emph{A bound for the maximum of a number of random
variables}, J. Appl. Prob. 3 (1966) 556--558.


\bibitem{Kounias}

E.~G. Kounias,  \emph{Bounds for the probability of a union, with
applications}, Ann. Math. Statist. 39 (1968) 2154--2158.

\bibitem{FengLiShen}
C. Feng, L. Li, J. Shen, \emph{On the Borel-Cantelli lemma and its
generalization}, C. R. Acad. Sci. Paris, Ser. I 347 (2009)
1313--1316.


\bibitem{FengLiShen2}
C. Feng, L. Li, J. Shen, \emph{Some inequalities in functional
analysis, combinatorics and probability theory}, Electronic J.
Combinatorics 17 (2010) \# R58.

\bibitem{Frolov}
A.~N. Frolov, \emph{Bounds for probabilities of unions of events and
the Borel-Cantelli lemma}, Statist. Probab. Lett. 82 (2012)
2189--2197.

\bibitem{Hoppe1}
F.~M. Hoppe, \emph{Improving probability bounds by optimization over
subsets}, Discrete Math. 306 (2006) 526--530.

\bibitem{Hoppe2}
F.~M. Hoppe, \emph{The effect of redundancy on probability bounds},
Discrete Math. 309 (2009) 123--127.

\bibitem{Hu}
S. Hu, X. Wang, X. Li, Y. Zhang, \emph{Comments on the paper: A
bilateral inequality on the Borel-Cantelli lemma},  Statist. Probab.
Lett. 79 (2009) 889--893.

\bibitem{KochenStone}
S. Kochen, C. Stone, \emph{A note on the Borel-Cantelli lemma},
Illnois J. Math. 8 (1964) 248--251.

\bibitem{KAT}
H. Kuai, F. Alajaji, G. Takahara, \emph{A lower bound on the
probability of a finite union of events}, Discrete Math. 215 (2000)
147--158.


\bibitem{Liu}
J. Liu, \emph{A note on the bilateral inequality for a sequence of
random variables},  Statist. Probab. Lett. 82 (2012) 871--875.

\bibitem{Mori}
T.~F. M\'{o}ri, G.~J. Sz\'{e}kely, \emph{On the Erd\H{o}s-R\'{e}nyi
generalization of the Borel-Cantelli lemma},  Studia Sci. Math.
Hungar.  18  (1983) 173--182.

\bibitem{Prekopa}
A. Pr\'{e}kopa, L. Gao, \emph{Bounding the probability of the union
of events by aggregation and disaggregation in linear programs},
Discrete Appl. Math. 145 (2005) 444--454.

\bibitem{Renyi}
A. R\'{e}nyi, Probability Theory, North-Holland Series in Applied
Mathematics and Mechanics, Vol. 10. North-Holland Publishing Co.,
Amsterdam-London, 1970; German version 1962, French version 1966,
new Hungarian edition 1965.


\bibitem{Spitzer}
F. Spitzer, Principles of Random Walk, Van Nostrand, Princeton,
1964.

%\bibitem{Xie}
%Y. Xie, \emph{A bilateral inequality on the Borel-Cantell lemma},
%Statist. Probab. Lett. 78 (2008) 2052--2057.

\bibitem{Xie}
Y. Xie, \emph{A bilateral inequality on a nonnegative bounded random
sequence}, Statist. Probab. Lett. 79 (2009) 1577--1580.



\end{thebibliography}

\documentclass[12pt,draft]{amsart}

\theoremstyle{definition}

\RequirePackage{amsmath}
\RequirePackage{amssymb}
\RequirePackage{amsthm}
 \RequirePackage{epsfig}

\begin{document}

\date{}

\title[On the Borel-Cantelli lemma]
{On the M\'{o}ri-Sz\'{e}kely conjectures for the Borel-Cantelli
lemma}

\author{Chunrong Feng and Liangpan Li}

\address{Department of Mathematical Sciences,
Loughborough University,  LE11 3TU,  UK}
 \email{C.Feng@lboro.ac.uk, L.Li@lboro.ac.uk}

\subjclass[2000]{60F15}

\keywords{Borel-Cantelli lemma, Gallot-Kounias bound,
Kuai-Alajaji-Takahara bound}

\date{}

\begin{abstract}
The purpose of this note is to show by constructing counterexamples
that two conjectures of M\'{o}ri and Sz\'{e}kely for the
Borel-Cantelli lemma are false.
\end{abstract}

\maketitle

\section{Introduction}
Let $\{A_n\}_{n=1}^{\infty}$ be an arbitrary sequence of events in a
probability space $(\Omega,\mathbb{P})$ and denote by $A_{\infty}$
the event that infinitely many $A_n$ occurs simultaneously, i.e.
\[A_{\infty}=\bigcap_{n=1}^{\infty}\bigcup_{k=n}^{\infty}A_k.\]
The classical Borel-Cantelli lemma states that:  if
$\sum_{n=1}^{\infty}\mathbb{P}(A_n)<\infty$ then
$\mathbb{P}(A_{\infty})=0$; else if
$\sum_{n=1}^{\infty}\mathbb{P}(A_n)=\infty$ and
$\{A_n\}_{n=1}^{\infty}$ are mutually independent, then
$\mathbb{P}(A_{\infty})=1$. In the past century many investigations
were devoted to the second implication in the attempt to weaken the
independence condition on $\{A_n\}_{n=1}^{\infty}$. For example, one
of the most applicable results is due to Erd\H{o}s and R\'{e}nyi
(\cite{ErdosRenyi,Renyi}, see also
\cite{Dawson,FengLiShen,KochenStone,Spitzer}) who proved that if
$\sum_{n=1}^{\infty}\mathbb{P}(A_n)=\infty$, then
\[
\mathbb{P}(A_{\infty})\geq\limsup_{n\rightarrow\infty}
\frac{\displaystyle\big(\sum_{k=1}^n{\mathbb{P}}(A_k)\big)^2}{\displaystyle\sum_{i=1}^n\sum_{j=1}^n{\mathbb{P}}(A_iA_j)}=
\limsup_{n\rightarrow\infty}\frac{\displaystyle1}{\displaystyle\mathbb{E}(\alpha_n^2)}\doteq\mbox{ER}(\{A_n\}),
\]
where $\mathbb{E}$ denotes the expectation function,
$\alpha_n\doteq{\displaystyle(\sum_{i=1}^n\mathbb{I}_{A_i})}/{\displaystyle(\sum_{i=1}^n\mathbb{P}({A_i}))}$,
and $\mathbb{I}_{A_i}$ is the indicator function of $A_i$. Later on,
by studying convex and concave Young functions M\'{o}ri and
Sz\'{e}kely (\cite{Mori}, see also
\cite{Amghibech,Frolov,Hu,Liu,Xie})
 improved the Erd\H{o}s-R\'{e}nyi bound to
\[
\mathbb{P}(A_{\infty})\geq\sup_{p\in(0,\infty)\backslash\{1\}}\big(
\limsup_{n\rightarrow\infty}(\mathbb{E}(\alpha_n^p)^{\frac{1}{1-p}})\big)=\lim_{p\searrow0}
\big(\limsup_{n\rightarrow\infty}\mathbb{E}(\alpha_n^p)\big)\doteq
\mbox{MS}(\{A_n\}).
\]
They also proposed the following two conjectures:

\textsc{Conjecture 1}:
\[\mathbb{P}(A_{\infty})=\sup\limits_{\tau:\mathbb{N}\rightarrow\mathbb{N}\ \mbox{is increasing}
}\mbox{MS}(\{A_{\tau(n)}\}).\]

\textsc{Conjecture 2}: If we have an estimate of the form
$\mathbb{P}(A_{\infty})\geq L_k$ $(k\geq2)$ where the constant $L_k$
depends only on $P(A_{i_1}A_{i_2}\cdots A_{i_k})$, $1\leq i_1 \leq
i_2\leq \cdots\leq i_k$, then
\[\sup\limits_{\tau:\mathbb{N}\rightarrow\mathbb{N}\ \mbox{is increasing}}\mbox{ER}(\{A_{\tau(n)}\})\geq L_k.\]

The purpose of this note is to show that both conjectures are false.

\section{A counterexample to Conjecture 1}

Let  $\mathbb{P}$ be the Lebesgue measure on $\Omega=[0,1]$ and let
\[A_{2^{i-1}+k}=[\frac{k}{2^i},\frac{k+1}{2^i}]\cup[\frac{1}{2},1]\ \
\ \ \ i\in\mathbb{N},\ 0\leq k<2^{i-1}.\] Obviously
$\mathbb{P}(A_{\infty})=1$. Let $p\in(0,1)$ and
$\tau:\mathbb{N}\rightarrow\mathbb{N}$ be any increasing function.
Then
\begin{align*}
\mathbb{E}\Big(\big(\frac{\displaystyle\sum_{i=1}^n\mathbb{I}_{A_{\tau(i)}}}{\displaystyle\sum_{i=1}^n\mathbb{P}({A_{\tau(i)}})}\big)^p\Big)&\leq
\frac{\displaystyle\mathbb{E}
\Big(\big({\sum_{i=1}^n\mathbb{I}_{A_{\tau(i)}}}\big)^p\Big)}{\displaystyle(\frac{n}{2})^p} \ \ \ \ \ \ \  \ \ \ \ \ \ \  \ \ \ \ \ \ \ \ \ \ \ \  \ \ \ \ \ \ \ \ \ (\mathbb{P}(A_{\tau(i)})\geq\frac{1}{2})\\
&\leq\frac{\displaystyle(\int_{[0,\frac{1}{2}]}\sum_{i=1}^n\mathbb{I}_{A_{\tau(i)}}d\mathbb{P})^p\cdot(\frac{1}{2})^{1-p}+\frac{n^p}{2}}
{\displaystyle(\frac{n}{2})^p} \ \ \ \ \ \ (\mbox{H\"{o}lder's inequality})\\
&\leq\frac{\displaystyle(\log_22n)^p\cdot(\frac{1}{2})^{1-p}+\frac{n^p}{2}}
{\displaystyle(\frac{n}{2})^p} \ \
 \big(\mathbb{P}(A_{\tau(i)}\cap[0,\frac{1}{2}])\leq\mathbb{P}(A_{i}\cap[0,\frac{1}{2}])\big).
\end{align*}
Letting first $n\rightarrow\infty$ then $p\rightarrow0$, we get
$\mbox{MS}(\{A_{\tau(n)}\})\leq\frac{1}{2}$. This example shows that
Conjecture 1 is false.

\section{A counterexample to Conjecture 2}

 In the beginning let us  recall two lower
bounds for  $\mathbb{P}(\cup_{i=1}^m A_i)$, where $\{A_i\}_{i=1}^m$
are finitely many events with non-zero probabilities in a
probability space $(\Omega,\mathbb{P})$. First, the Gallot-Kounias
bound (\cite{Gallot,Kounias}, see also \cite{FengLiShen2,Frolov} for
more details) claims that
\begin{equation}\label{GK}
\mathbb{P}(\cup_{i=1}^mA_i)\geq\max_{(\omega_1,\ldots,\omega_m)\in\mathbb{R}^m}
 \frac{\displaystyle\big(\sum_{i=1}^m\omega_i \mathbb{P}(A_i)\big)^2}{\displaystyle\sum_{i=1}^m\sum_{j=1}^m\omega_i\omega_j \mathbb{P}(A_{i}A_j)}
 =\sum_{i=1}^m\gamma_i\doteq \mbox{GK}(\{A_i\}_{i=1}^m),
\end{equation}
 where $\frac{0}{0}\doteq0$ and $(\gamma_1,\ldots,\gamma_m)\in\mathbb{R}^m$ is any solution to
\begin{equation}\label{GK2}
 \Big(\frac{\mathbb{P}(A_{i}A_j)}{\mathbb{P}(A_{i})\mathbb{P}(A_j)}\Big)_{m\times m} \left(
\begin{array}{c}
\gamma_1 \\
\vdots\\
\gamma_m
\end{array} \right )=\left ( \begin{array}{c}
1 \\
\vdots\\
1
\end{array} \right).
\end{equation} Second, Kuai, Alajaji and Takahara (\cite{KAT}, see
also \cite{Hoppe1,Hoppe2,Prekopa}) proved that
\begin{equation}
\label{KAT} \mathbb{P}(\cup_{i=1}^mA_i) \geq \sum _{i=1}^m \Big(
\frac {\displaystyle\theta_i \mathbb{P}(A_i)^2}{\displaystyle S_i+
(1-\theta_i)\mathbb{P}(A_i)}+ \frac
{\displaystyle(1-\theta_i)\mathbb{P}(A_i)^2}{\displaystyle
S_i-\theta_i\mathbb{P}(A_i)} \Big)\doteq
\mbox{KAT}(\{A_i\}_{i=1}^m), \end{equation} where
$S_i\doteq\sum_{j=1}^m\mathbb{P}(A_iA_j)$, $\theta_i$ is the
fractional part of $\frac{S_i}{\mathbb{P}(A_i)}$.

Next let us explain how will we find a  counterexample to Conjecture
2. Suppose  a sequence of events $\{A_n\}_{n=1}^{\infty}$ with
non-zero probabilities occur periodically as follows:
\[A_1,A_2,\ldots,A_m,A_1,A_2,\ldots,A_m,A_1,A_2,\ldots,A_m,\ldots.\]
Obviously,
\[\mathbb{P}(A_{\infty})
\geq\limsup_{n\rightarrow\infty}\mathbb{P}(\bigcup_{i=n}^{n+m-1}A_i)\geq\limsup_{n\rightarrow\infty}\mbox{KAT}(\{A_i\}_{i=n}^{n+m-1})=\mbox{KAT}(\{A_i\}_{i=1}^m).\]
On the other hand, it is easy to observe that
\[\sup\limits_{\tau:\mathbb{N}\rightarrow\mathbb{N}\ \mbox{is increasing}}\mbox{ER}(\{A_{\tau(n)}\})\leq\max_{(\omega_1,\ldots,\omega_m)\in\mathbb{R}^m}
 \frac{\displaystyle\big(\sum_{i=1}^m\omega_i \mathbb{P}(A_i)\big)^2}{\displaystyle\sum_{i=1}^m\sum_{j=1}^m\omega_i\omega_j \mathbb{P}(A_{i}A_j)}
 =\mbox{GK}(\{A_i\}_{i=1}^m).
 \]
Hence to disprove Conjecture 2 it suffices to construct finitely
 many
 $\{A_i\}_{i=1}^m$ so that
 $\mbox{GK}(\{A_i\}_{i=1}^m)<\mbox{KAT}(\{A_i\}_{i=1}^m)$.

  To this aim consider six events $\{A_i\}_{i=1}^6$ in a finite probability space
\begin{eqnarray*}
&\begin{tabular}{|c|c|c|c|c|c|c|c|c|}
\hline $x$ & $\mathbb{P}(\{x\})$ & $A_1$ & $A_2$ & $A_3$ & $A_4$ & $A_5$ & $A_6$\\
\hline $x_1$ & 0.2 & $\bigstar$ & $\bigstar$ & $\bigstar$ & & & $\bigstar$   \\
\hline $x_2$       & 0.2 & $\bigstar$ & & &$\bigstar$ &$\bigstar$ &$\bigstar$ \\
\hline $x_3$       & 0.2 & &$\bigstar$ & $\bigstar$ & & & $\bigstar$ \\
\hline $x_4$       & 0.2 &$\bigstar$ & & & $\bigstar$ & $\bigstar$ &\\
\hline $x_5$       & 0.2 & &$\bigstar$ & $\bigstar$ &$\bigstar$ &$\bigstar$ &  \\
\hline
\end{tabular}
\end{eqnarray*}
with joint probability matrix
\[
(\mathbb{P}(A_iA_j))=\left ( \begin{array}{cccccc}
  0.6  & 0.2 & 0.2 & 0.4 & 0.4& 0.4\\
  0.2 & 0.6 & 0.6 & 0.2 & 0.2& 0.4\\
 0.2 & 0.6 &0.6 & 0.2 & 0.2& 0.4 \\
0.4& 0.2 &0.2 &0.6 & 0.6& 0.2 \\
0.4& 0.2 &0.2 &0.6 &0.6 &0.2\\
0.4& 0.4 &0.4 &0.2 & 0.2&0.6
\end{array} \right ).
\]
Then it is straightforward to work out from
(\ref{GK})$\sim$(\ref{KAT}) that
\[\mbox{GK}(\{A_i\}_{i=1}^6)=\frac{54}{55}<1=\mbox{KAT}(\{A_i\}_{i=1}^6).\]
We are done.


\end{document}